Babel, a multilingual package for use with L^{A}TEX’s standard document classes^{∗}

Johannes Braams Kooienswater 62 2715 AJ Zoetermeer The Netherlands

JLBraams@cistron.nl

Printed April 11, 1998

Abstract

The standard distribution of L^{A}TEX contains a number of document classes that are meant to be used, but also serve as examples for other users to create their own document classes. These document classes have become very popular among L^{A}TEX users. But it should be kept in mind that they were designed for American tastes and typography. At one time they contained a number of hard-wired texts. This report describes babel, a package that makes use of the new capabilities of TEX version 3 to provide an environment in which documents can be typeset in a non-american language or in more than one language.

Contents

1 The user interface 2

1.1 Languages supported by

Babel . . . . . . . . . . . . 3 1.2 Workarounds . . . . . . . 4

2 Changes for L^{A}TEX 2ε 5

3 Changes in Babel version 3.6 5

4 Changes in Babel version 3.5 6

5 The interface between the core of babel and the language definition files 7 5.1 Support for active char-

acters . . . . . . . . . . . 8 5.2 Support for saving macro

definitions . . . . . . . . . 9 5.3 Support for extending

macros . . . . . . . . . . . 9 5.4 Macros common to a

number of languages … 9

6 Compatibility with german.sty 9

7 Compatibility with the french package 10

8 Conclusion 11

9 Acknowledgements 11

10 The Esperanto language 12

11 The Dutch language 12 12 The English language 12

∗During the development ideas from Nico Poppelier, Piet van Oostrum and many others have been used. Bernd Raichle has provided many helpful suggestions.

13 The German language 13

14 The Breton language 13

15 The Welsh language 13

16 The Irish language 14

17 The Scottish language 15

18 The Greek language 15

19 The French Language 15 19.1 About French typography 15

20 The Italian language 18

21 The Portuguese language 18

22 The Spanish language 18

23 The Catalan language 20

24 The Galician language 21

25 The Romanian language 21

26 The Danish language 22 27 The Norwegian language 22

28 The Swedish language 22

29 The Finnish language 22

30 The Hungarian language 23

31 The Estonian language 24

32 The Croatian language 25

33 The Czech language 25

34 The Polish language 25

35 The Slovak language 25

36 The Slovenian language 25

37 The Russian language 26

38 The Lower Sorbian language 27

39 The Upper Sorbian language 27

40 The Turkish language 28

41 The Bahasa language 28

1 The user interface

The user interface of this package is quite simple. It consists of a set of commands that switch from one language to another and a set of commands that deal with shorthands. It is also possible to find out out what the current language is.

\selectlanguage When a user wants to switch from one language to another he can do so using the macro \selectlanguage. This macro takes the language, defined previously by a language definition file, as its argument. It calls several macros that should be defined in the language definition files to activate the special definitions for the language chosen.

otherlanguage The environment otherlanguage does basically the same as \selectlanguage, except the language change is local to the environment. For mixing left-to-right typesetting with right-to-left typesetting the use of this environment is a prerequisite. The language to switch to is specified as an argument to

\begin{otherlanguage}.

\foreignlanguage The command \foreignlanguage takes two arguments, the second argument is a phrase to be typeset according to the rules of the language named in its first argument. This command only switches the extra definitions and the hyphenation rules for the language, not the names and dates. otherlanguage* In the environment otherlanguage* only the typesetting is done according to

the rules of the other language, but the text-strings such as ‘figure’, ‘table’, etc. are left as they were set outside this environment.

\languagename The control sequence \languagename contains the name of the current lan-

guage.

\iflanguage If more than one language is used it might be necessary to know which language is active at a specific time. This can be checked by a call to \iflanguage. This macro takes three arguments. The first argument is the name of a language, the second and third arguments are the actions to take if the result of the test is true or false respectively.

\useshorthands The command \useshorthands initiates the definition of user-defined shorthand sequences. It has one argument, the character which starts these personal shorthands.

\defineshorthand The command \defineshorthand takes two arguments, the first of which is a one or two character sequence, the second argument is the code the shorthand should expand to.

\aliasshorthand The command \aliasshorthand can be used to let another character perform the same functions as the default shorthand character. If one prefers for example to use the character / over " in typing polish texts this can be acheived by entering

\aliasshorthand{"}{/}.

\languageshorthands The command \languageshorthands can be used to switch the shorthands on the language level. It takes one argument, the name of a language. Note that for this to work the language should have been specified as an option when loading the babel package.

1.1 Languages supported by Babel

In the following table all the languages supported by Babel are listed, together with the names of the options with which you can load babel for each language.

Afrikaans  afrikaans
Bahasa  bahasa
Breton  breton
Catalan  catalan
Croatian  croatian
Czech  czech
Danish  danish
Dutch  dutch
English  english, USenglish, american, UKenglish, british
Esperanto  esperanto
Estonian  estonian
Finnish  finnish
French  french, francais
Galician  galician
German  austrian, german, germanb
Greek  greek
Hebrew  hebrew
Hungarian  magyar, hungarian
Irish Gaelic  irish
Italian  italian
Lower Sorbian lowersorbian
Norwegian  norsk, nynorsk
Polish  polish
Portuguese  portuges, portuguese, brazilian, brazil
Romanian  romanian
Russian  russian
Scottish Gaelic scottish
Spanish  spanish
Slovakian  slovak
Slovenian  slovene
Swedish  swedish
Turkish  turkish
Upper Sorbian uppersorbian
Welsh  welsh

For some languages babel supports the options activeacute and activegrave; for typestting russian texts babel knows about the options LWN and LCY to specify the fontencoding of the cyrillic font used. Currently only LWN is supported.

1.2 Workarounds

When you use the document class book and you use \ref inside the argument of \chapter you will experience the problem that L^{A}TEX will keep complaining about an undefined label. The reason is that the argument of \ref is passed through \uppercase at some time during processing. To prevent such problems you could revert to using uppercase labels, or you can use \lowercase{\ref{foo}} inside the argument of \chapter.

2 Changes for L^{A}TEX 2ε

With the advent of L^{A}TEX 2ε the interface to babel in the preamble of the doument has changed. With L^{A}TEX2.09 one used to call up the babel system with a line such as:

\documentstyle[dutch,english]{article}

which would tell L^{A}TEX that the document would be written in two languages, dutch and english and that english would be the first language in use.

The L^{A}TEX 2ε way of providing the same information is:

\documentclass{article}
\usepackage[dutch,english]{babel}

or, making dutch and english global options in order to let other packages detect and use them:

\documentclass[dutch,english]{article}
\usepackage{babel}
\usepackage{varioref}

In this last example the package varioref will also see the options and will be able to use them.

3 Changes in Babel version 3.6

In Babel version 3.6 a number of bugs that were found in version 3.5 are fixed. Also a number of changes and additions have occured:

• A new environment otherlanguage* is introduced. it only switches the ‘specials’, but leaves the ‘captions’ untouched.

• The shorthands are no longer fully expandable. Some problems could only be solved by peeking at the token following an active character. The advantage is that '{}a works as expected for languages that have the ' active.

• Support for typesetting french texts is much enhanced; the file francais.ldf is now replaced by frenchb.ldf which is maintained by Daniel Flipo.

• Support for typesetting the russian language is again available. The languange definition file was originally developped by Olga Lapko from cyrtug. The fonts needed to typeset the russian language are now part of the babel distribution. The support is not yet up to the level which is needed according to Olga, but this is a start.

• Support for typesetting greek texts is now also available. What is offered in this release is a first attempt; it will be enhanced later on by Yannis Haralambous.

• in babel 3.6j some hooks have been added for the development of support for Hebrew typesetting.

• Support for typesetting texts in Afrikaans (a variant of Dutch, spoken in South Africa) has been added to dutch.ldf.

• Support for typesetting welsh texts is now available.

• A new command \aliasshorthand is introduced. It seems that in Poland various conventions are used to type the necessary polish letters. It is now possible to use the character / as a shorthand character instead of the character " by issuing the command \aliasshorthand{"}{/}.

• The shorthand mechanism now deals correctly with characters that are already active.

• Shorthand characters are made active at \begin{document}, not earlier. This is to prevent problems with other packages.

• A preambleonly command \substitutefontfamily has been added to create .fd files on the fly when the font families of the latin text differ from the families used for the cyrillic or greek parts of the text.

• Three new comands \LdfInit, \ldf@quit and \ldf@finish are introduced that perform a number of standard tasks.

4 Changes in Babel version 3.5

In Babel version 3.5 a lot of changes have been made when compared with the previous release. Here is a list of the most important ones:

• the selection of the language is delayed untill \begin{document}, this has the consequence that you need to add appropriate \selectlanguage commands if you include \hyphenation lists in the preamble of your document.

• babel now has a language environment and a new command \foreignlanguage;

• the way active characters are dealt with is completely changed. They are called ‘shorthands’; one can have three levels of shorthands: on the user level, the language level and on ‘system level’. A consequence of the new way of handling active characters is that they are now written to auxiliary files ‘verbatim’;

• A language change now also writes information in the .aux file as the change might also affect typesetting the table of contents. The consequence is that an .aux file generated by a LaTeX format with babel preloaded gives errors when read with a LaTeX format without babel, but I think this problaly doesn’t occur;

• babel is now compatible with the inputenc and fontenc packages;

• the language definition files now have a new extension, ldf;

• the syntax of the file language.dat is extended to be compatible with the french package by Bernard Gaulle;

• each language definition file looks for a configuration file which has the same name, but the extension .cfg. It can cantain any valid L^{A}TEX code.

5 The interface between the core of babel and the language definition files

In the core of the babel system two macros are defined that are to be used in language definition files. Their purpose is to make a new language known.

\addlanguage The macro \addlanguage is a non-outer version of the macro \newlanguage, defined in plain.tex version 3.x. For older versions of plain.tex and lplain.tex a substitute definition is used.

\adddialect The macro \adddialect can be used in the case where two languages can (or have to) use the same hyphenation patterns. This can alos be useful when a user wants to use a language for which no patterns are preloaded in the format. In such a case the default behaviour of the babel system is to define this language as a ‘dialect’ of the language for which the patterns were loaded as \language0.

The language definition files have to conform to a number of conventions. The reason for this is that these files have to fill in the gaps left by the common code in babel.def, i. e., the definitions of the macros that produce texts. Also the language-switching possibility which has been built into the babel system has its implications.

The following assumptions are made:

• Some of the language-specific definitions might be used by plain TEX users, so the files have to be coded such that they can be read by L^{A}TEX as well as by plain TEX. The current format can be checked by looking at the value of the macro \fmtname.

• The common part of the babel system redefines a number of macros and environments (defined previously in the document style) to put in the names of macros that replace the previously hard-wired texts. These macros have to be defined in the language definition files.

• The language definition files define five macros, used to activate and deactivate the language-specific definitions. These macros are \⟨lang⟩hyphenmins, \captions⟨lang⟩, \date⟨lang⟩, \extras⟨lang⟩and \noextras⟨lang⟩; where ⟨lang⟩ is either the name of the language definition file or the name of the L^{A}TEX option that is to be used. These macros and their functions are discussed below.

• When a language definition file is loaded, it can define \l@⟨lang⟩to be a dialect of \language0 when \l@⟨lang⟩is undefined.

• The languagedefinition files can be read in the preamble of the document, but also in the middle of document processing. This means that they have to function independently of the current \catcode of the @ sign.

\langhyphenmins The macro \⟨lang⟩hyphenmins is used to store the values of the \lefthyphenmin

and \righthyphenmin.

\captionslang The macro \captions⟨lang⟩defines the macros that hold the texts to replace

the original hard-wired texts.

\datelang The macro \date⟨lang⟩defines \today and \extraslang The macro \extras⟨lang⟩contains all the extra definitions needed for a specific

language.

\noextraslang Because we want to offer the user the possibility to switch between languages and we do not know in what state TEX might be after the execution of \extras⟨lang⟩, a macro that brings TEX into a predefined state is needed. It will be no surprise that the name of this macro is \noextras⟨lang⟩.

\main@language To postpone the activation of the definitions needed for a language untill the beginning of a document, all language definition files should use \main@language instead of \selectlanguage. This will just store the name of the language and the proper language will be activated at the start of the document.

\LdfInit The macro \LdfInit performs a couple of standard checks that have to be made at the beginning of a language definition file, such as checking the category code of the @-sign, preventing that the .ldf file is processed twice, etc.

\ldf@quit The macro \ldf@quit performs a couple of tasks that need to be taken care of when a .ldf file was processed earlier. These tasks include the resetting of the category code of the @-sign, preparing the language to be activated at \begin{document} time and ending the input stream.

\ldf@finish The macro \ldf@finish performs a couple of tasks that need to be taken care of at the end of each .ldf file. These tasks include the resetting of the category code of the @-sign, the loading of a local configuration file and preparing the language to be activated at \begin{document} time.

\loadlocalcfg At the end of the processing of a language definition file L^{A}TEX can be instructed to load a local configuration file. This file can for instance be used to add strings to \captions⟨lang⟩ in order to support local document classes. The user will be informed of the fact that this configuration file is loaded. This macro is called by

\finish@ldf.

\subsitutefontfamily This command takes three arguments, a font encoding and two font family names. It creates a font description file for the first font in the given encoding. This .fd file will instruct L^{A}TEX to use a font from the second family when a font from the first family in the given encoding seems to needed.

5.1 Support for active characters

In quite a number of language definition files, active characters are introduced. To facilitate this, some support macros are provided.

\initiate@active@char The internal macro \initiate@active@char is used in language definition files to instruct L^{A}TEX to give a character the category code ‘active’. When a character has been made active it will remain that way untill the end of then document. Its definition may vary.

\bbl@activate The command \bbl@activate is used to change the way an active character \bbl@deactivate expands. \bbl@activate ‘switches on’ the active behaviour of the character.

\bbl@deactive lets the active character expand to its former (mostly) non-active self.

\declare@shorthand The macro \declare@shorthand is used to define the various shorthands. It takes three arguments, the name for the collection of shorthands this definition belongs to; the character (sequence) that makes up the shorthand i.i. ~ or "a and the code to be executed when the shorthand is encountered.

\bbl@add@special The TEXbook states: “Plain TEX includes a macro called \dospecials that \bbl@remove@special is essentially a set macro, representing the set of all characters that have a spe-

cial category code.” [1, p. 380] It is used to set text ‘verbatim’. To make this work if more characters get a special category code, you have to add this character to the macro \dospecial. L^{A}TEX adds another macro called \@sanitize representing the same character set, but without the curly braces. The macros

\bbl@add@special⟨char⟩ and \bbl@remove@special⟨char⟩ add and remove the

character ⟨char⟩ to these two sets.

5.2 Support for saving macro definitions

Language definition files may want to redefine macros that already exist. Therefore a mechanism for saving (and restoring) the original definition of those macros is provided. We provide two macros for this^{1}.

\babel@save To save the current meaning of any control sequence the macro \babel@save is provided. It takes one argument, ⟨csname⟩, the control sequence for which the meaning has to be saved.

\babel@savevariable A second macro is provided to save the current value of a variable. In this context anything that is allowed after the \the primitive is considered to be a variable. The macro takes one argument, the ⟨variable⟩.

The effect of the aforementioned macros is that a piece of code is appended to the current definition of \originalTeX. When \originalTeX is expanded this code restores the previous definition of the control sequence or the previous value of the variable.

5.3 Support for extending macros

\addto The macro \addto{⟨control sequence⟩}{⟨TEX code⟩} can be used to extend the definition of a macro. The macro need not be defined. This macro can, for instance, be used in adding instructions to a macro like \extrasenglish.

5.4 Macros common to a number of languages

\allowhyphens In a couple of european languages compound words are used. This means that when TEX has to hyphenate such a compound word it only does that at the ‘-’ that is used in such words. To allow hyphenation in the rest of such a compound word the macro \allowhyphens can be used.

\set@low@box For some languages quotes need to be lowered to the baseline. For this purpose the macro \set@low@box is available. It takes one argument and puts that argument in an \hbox, at the baseline. The result is available in \box0 for further processing.

\save@sf@q Sometimes it is necessary to preserve the \spacefactor. For this purpose the macro \save@sf@q is available. It takes one argument, saves the current spacefactor, executes the argument and restores the spacefactor.

\bbl@frenchspacing  The commands \bbl@frenchspacing and \bbl@nonfrenchspacing can be

\bbl@nonfrenchspacing used to properly switch french spacing on and off.

6 Compatibility with german.sty

The file german.sty has been one of the sources of inspiration for the babel system. Because of this I wanted to include german.sty in the babel system. To be able to do that I had to allow for one incompatibility: in the definition of the macro \selectlanguage in german.sty the argument is used as the ⟨number⟩ for an \ifcase. So in this case a call to \selectlanguage might look like

1This mechanism was introduced by Bernd Raichle.

\selectlanguage{\german}.

In the definition of the macro \selectlanguage in babel.def the argument is used as a part of other macronames, so a call to \selectlanguage now looks like \selectlanguage{german}. Notice the absence of the escape character. As of version 3.1a of babel both syntaxes are allowed.

All other features of the original german.sty have been copied into a new file,

called germanb.sty^{2}.

Although the babel system was developed to be used with L^{A}TEX, some of the features implemented in the language definition files might be needed by plain TEX users. Care has been taken that all files in the system can be processed by plain TEX.

7 Compatibility with the french package

It has been reported to me that the package french by Bernard Gaulle (gaulle@idris.fr) works together with babel.

Therefore, babel will first search for the file french.ldf when you give it the option french; then it will try to load frenchb.ldf. When you give babel the option francais it will only look for frenchb.ldf.

2The ‘b’ is added to the name to distinguish the file from Partls’ file.

8 Conclusion

A system of document options has been presented that enable the user of L^{A}TEX to adapt the standard document classes of L^{A}TEX to the language he or she prefers to use. These options offer the possibility to switch between languages in one document. The basic interface consists of using ones option, which is the same for all standard document classes.

In some cases the language definition files provide macros that can be of use to plain TEX users as well as to L^{A}TEX users. The babel system has been implemented in such a way that it can be used by both groups of users.

9 Acknowledgements

I would like to thank all who volunteered as β-testers for their time. I would like to mention Julio Sanchez who supplied the option file for the Spanish language and Maurizio Codogno who supplied the option file for the Italian language. Michel Goossens supplied contributions for most of the other languages. Nico Poppelier helped polishing the text of the documentation and supplied parts of the macros for the Dutch language. Paul Wackers and Werenfried Spit helped finding and repairing bugs.

During the further development of the babel system I received much help from Bernd Raichle, for which I am grateful.

References

[1] Donald E. Knuth, The TEXbook, Addison-Wesley, 1986.

[2] Leslie Lamport, L^{A}TEX, A document preparation System, Addison-Wesley, 1986.

[3] K.F. Treebus. Tekstwijzer, een gids voor het grafisch verwerken van tekst. SDU Uitgeverij (’s-Gravenhage, 1988). A Dutch book on layout design and typography.

[4] Hubert Partl, German TEX, TUGboat 9 (1988) #1, p. 70–72.

[5] Leslie Lamport, in: TEXhax Digest, Volume 89, #13, 17 februari 1989.

[6] Johannes Braams, Victor Eijkhout and Nico Poppelier, The development of national L^{A}TEX styles, TUGboat 10 (1989) #3, p. 401–406.

[7] Joachim Schrod, International L^{A}TEX is ready to use, TUGboat 11 (1990) #1, p. 87–90.

10 The Esperanto language

The file esperant.dtx^{3} defines all the language-specific macros for the Esperanto language.

For this language the character ^is made active. In table 1 an overview is given of its purpose.

^c gives ĉ with hyphenation in the rest of the word allowed, this works for c, C, g, G, H, J, s, S, z, Z

^h prevents ĥ from becoming too tall
^j gives ĵ

^u gives ŭ, with hyphenation in the rest of the word

allowed ^U gives Ŭ, with hyphenation in the rest of the word

allowed

          ^| inserts a \discretionary{-}{}{}

Table 1: The functions of the active character for Esperanto.

11 The Dutch language

The file dutch.dtx^{4} defines all the language-specific macros for the Dutch language and the ‘Afrikaans’ version^{5} of it.

For this language the character " is made active. In table 2 an overview is given of its purpose. One of the reasons for this is that in the Dutch language a word with a dieresis can be hyphenated just before the letter with the umlaut, but the dieresis has to disappear if the word is broken between the previous letter and the accented letter.

In [3] the quoting conventions for the Dutch language are discussed. The preferred convention is the single-quote Anglo-American convention, i.e. ‘This is a quote’. An alternative is the slightly old-fashioned Dutch method with initial double quotes lowered to the baseline, ”This is a quote”, which should be typed

as "`This is a quote"'.

12 The English language

The file english.dtx^{6} defines all the language definition macros for the English language as well as for the American version of this language.

For this language currently no special definitions are needed or available.

3The file described in this section has version number v1.4j and was last revised on 1997/01/06.

A contribution was made by Ruiz-Altaba Marti (ruizaltb@cernvm.cern.ch). Code from the file esperant.sty by Jörg Knappen (knappen@vkpmzd.kph.uni-mainz.de) was included.

4The file described in this section has version number v3.8c, and was last revised on 1997/01/14.

5contributed by Stoffel Lombard (lombc@b31pc87.up.ac.za)

6The file described in this section has version number v3.3h and was last revised on 1996/12/23.

"a \"a which hyphenates as -a; also implemented for

the other letters.

"y puts a negative kern between i and j
"Y puts a negative kern between I and J
"| disable ligature at this position.

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word. "~ to produce a hyphencharcter without the following

              \discretionary{}{}{}.

"" to produce an invisible ‘breakpoint’.
"` lowered double left quotes (see example below).
"' normal double right quotes.

\- like the old \-, but allowing hyphenation in the rest of the word.

Table 2: The extra definitions made by dutch.ldf

13 The German language

The file germanb.dtx^{7} defines all the language definition macros for the German language as well as for the Austrian dialect of this language^{8}.

For this language the character " is made active. In table 3 an overview is given of its purpose. One of the reasons for this is that in the German language some character combinations change when a word is broken between the combination. Also the vertical placement of the umlaut can be controlled this way. The quotes in table 3 can also be typeset by using the commands in table 4.

14 The Breton language

The file breton.dtx^{9} defines all the language-specific macros for the Breton language.

There are not really typographic rules for the Breton language. It is a local language (it’s one of the celtic languages) which is spoken in Brittany (West of France). So we have a synthesis between french typographic rules and english typographic rules. The characters :, ;, ! and ? are made active in order to get a whitespace automatically before these characters.

15 The Welsh language

The file welsh.dtx^{10} defines all the language definition macros for the Welsh language as well as for the ¡Dialect¿ version of this language.

For this language currently no special definitions are needed or available.

7The file described in this section has version number v2.6d and was last revised on 1996/12/23.

8This file is a re-implementation of Hubert Partl’s german.sty version 2.5b, see [4].

9The file described in this section has version number v1.0e and was last revised on 1996/12/23.

10The file described in this section has version number v1.0a and was last revised on 1996/12/23.

"a \"a, also implemented for the other lowercase and

uppercase vowels.

"s to produce the German ß (like \ss{}).
"z to produce the German ß (like \ss{}).
"ck for ck to be hyphenated as k-k.

"ff for ff to be hyphenated as ff-f, this is also imple-

mented for l, m, n, p, r and t

"S for SS to be \uppercase{"s}.
"Z for SZ to be \uppercase{"z}.

"| disable ligature at this position.

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word. "" like "-, but producing no hyphen sign (for compund

words with hyphen, e.g. x-""y).

"~ for a compound word mark without a breakpoint.

"= for a compound word mark with a breakpoint, allow-

ing hyphenation in the composing words.

"` for German left double quotes (looks like ,,).
"' for German right double quotes.
"< for French left double quotes (similar to <<).
"> for French right double quotes (similar to >>).

Table 3: The extra definitions made by german.ldf

16 The Irish language

The file irish.dtx^{11} defines all the language definition macros for the Irish language.

For this language currently no special definitions are needed or available.

11The file described in this section has version number v1.0e and was last revised on 1996/12/23. A contribution was made by Marion Gunn.

\glqq for German left double quotes (looks like ,,).
\grqq for German right double quotes (looks like “).
\glq for German left single quotes (looks like ,).
\grq for German right single quotes (looks like ‘).
\flqq for French left double quotes (similar to <<).
\frqq for French right double quotes (similar to >>).
\flq for (French) left single quotes (similar to <).
\frq for (French) right single quotes (similar to >).
\dq  the original quotes character (").

Table 4: More commands which produce quotes, defined by german.ldf

17 The Scottish language

The file scottish.dtx^{12} defines all the language definition macros for the Scottish language.

For this language currently no special definitions are needed or available.

18 The Greek language

The file greek.dtx^{13} defines all the language definition macros for the Greek language.

19 The French Language

19.1 About French typography

The file frenchb.dtx^{14}, derived from frenchy.sty, defines all the language definition macros for the French language.

Customization for the French language is achieved following the book “Lexique des règles typographiques en usage à l’Imprimerie nationale” troisième édition (1994), ISBN-2-11-081075-0.

This file has been designed to be used with L^{A}TEX 2ε, L^{A}TEX-2.09 and PlainTEX formats. If you are still using L^{A}TEX-2.09, you should consider switching to L^{A}TEX 2ε!

   Any of the commands \selectlanguage{french}, \selectlanguage{francais},

or \selectlanguage{frenchb} switches to the French language with the following effects:

1. French hyphenation patterns are made active;

2. ‘double punctuation’ is made active for correct spacing in French;

3. \today prints the date in French;

4. the caption names are translated into French (L^{A}TEX only);

5. the list items are set to ‘–’ instead of • (L^{A}TEX only);

6. the vertical spacing in lists is shortened (L^{A}TEX only);

7. the first paragraph of each section is indented (L^{A}TEX only);

8. French quotation marks can be typeset using the commands \og and \fg which work in L^{A}TEX 2ε, L^{A}TEX-2.09 and PlainTEX, their appearance depending on what is available to draw them; if you use L^{A}TEX 2ε with T1-encoding you can also enter them as <<~French quotation marks~>> but then don’t forget the unbreakable spaces, (\og and \fg provide for correct line breaks); 9. a command \up is provided to typeset superscripts like M\up{me} (abbrevi-

12The file described in this section has version number v1.0d and was last revised on 1996/12/23. A contribution was made by Fraser Grant (FRASER@CERNVM).

13The file described in this section has version number v1.0b and was last revised on 1996/12/23. The original author is Apostolos Syropoulos (apostolo@platon.ee.duth.gr).

14The file described in this section has version number v1.2a and was last revised on 1997/01/11.

ation for “Madame”), 1\up{er} (for “premier”);

10. family names should be typeset in small capitals and never be hyphenated, the macro \bsc (boxed small caps) does this, e.g., Leslie~\bsc{Lamport} will produce Leslie Lamport;

11. commands \primo, \secundo, \tertio and \quarto may be used to enumerate in lists;

12. abbreviations for “Numéro” and “numéro” are obtained via the commands

     \No, \no;

13. two commands are provided to typeset abbreviations for “degré”: \degre prints the raw character and \degres should be used to typeset temperatures (e.g., “20~\degres C” with an unbreakable space), or for alcohols’ strengths (e.g., “45\degres” with no space in French);

14. an command \nombre is provided to ease the typesetting of numbers: it works both in text and in math-mode: inputting \nombre{3141,592653} will format this number properly according to the current language (French or non-French)^{15}. The command \nombre is a contribution of Vincent Jalby using ideas of David Carlisle in comma.sty.

All commands previously available in francais.ldf have been included in frenchb.ldf for compatibility, sometimes with updated definitions.

The french package, by Bernard Gaulle, was not designed to run with babel (although the latest versions claim to be babel compatible), but rather as a standalone package for the French language. It provides many more functionalities (like \lettrine, \sommaire…) not available in frenchb, at the cost of a much greater complexity and possible incompatibilities with other languages.

As french is known to produce the best layout available for French typography, I have borrowed many ideas from Bernard’s file. I did my best to help users of both packages (french and frenchb) to exchange their sources files easily, with one exception which affects the way French quotation marks are entered: frenchb uses macros (\og and \fg) while french uses active characters (<< and >>).

French typographic rules specify that some white space should be present before ‘double punctuation’ characters. These characters are ;!? and :. In order to get this white space automatically, the category code of these characters is made \active. In French, the user should input these four characters preceded with a space, but as many people forget about it (even among native French writers!), the default behaviour of frenchb is to automatically add a \thinspace before ‘;’ ‘!’ ‘?’ and a normal (unbreakable) space before ‘:’ (this is the rule in French typography). It’s up to the user to add or not a space after ‘double punctuation’ characters: usually a space is necessary, but not always (before a full point or a closing brace for instance), so this cannot done automatically.

15In math-mode the comma has to be surrounded with braces to avoid a spurious space being inserted after it (see the TEXbook p. 134). Besides this, each slice of three digits should be separated either with a comma in English or with a space in French.

In (rare) cases where no space should be added before a ‘double punctua-

tion’, either use \string; \string: \string! \string? instead of ; : ! ?,

or switch locally to english. For instance you can type C\string:TEX or

\begin{otherlanguage}{english}{C:TEX}\end{otherlanguage} to avoid the

space before : in a MS-DOS path.

Some users dislike this automatic insertion of a space before ‘double punctuation’, and prefer to decide themselves whether a space should be added or not; so a hook \NoAutoSpaceBeforeFDP is provided: if this command is added (in file frenchb.cfg, or anywhere in a document) frenchb will respect your typing, and introduce a suitable space before ‘double punctuation’ if and only if a space is typed in the source file before those signs.

The command \AutoSpaceBeforeFDP switches back to the default behavior of

frenchb.

Once you have built your format, a good precaution would be to perform some basic tests about hyphenation in French. For L^{A}TEX 2ε I suggest this:

• run the following file, with the encoding suitable for your machine (myencoding will be latin1 for unix machines and PCs running Windows, applemac for Macintoshs, or cp850 for PCs running DOS. If you are using MlTEX together with CMR fonts, comment out the line

\usepackage[my-encoding]{inputenc}.
%%% Test file for French hyphenation.
\documentclass{article}
\usepackage[my-encoding]{inputenc}
\usepackage[francais]{babel}
\begin{document}
\showhyphens{signal, container, \'ev\'enement, alg\`ebre}
\showhyphens{signal, container,événement, algèbre}
\end{document}

• check the hyphenations proposed by TEX in your log-file; in French you should get with both 7-bit and 8-bit encodings

si-gnal, contai-ner, évé-ne-ment, al-gèbre.

Do not care about how accented characters are displayed in the log-file, what matters is the position of the ‘-’ hyphen signs only.

If they are all correct, your installation (probably) works fine, if one (or more) is (are) wrong, ask a local wizard to see what’s going wrong and perform the test again (or e-mail me about what happens). Frequent mismatches:

• you get sig-nal, con-tainer, this probably means that the hyphenation patterns you are using are for USenglish, not for French;

• you get no hyphen at all in évé-ne-ment, this probably means that you are using CMR fonts and the macro \accent to produce accented characters. Consider switching to DC/EC fonts and T1-encoding or use MlTEX.

frenchb has been improved using helpful suggestions from many people, the main contributions came from Vincent Jalby. Thanks to all of them!

First version released: 1.1 as of 1996/05/31 part of babel-3.6beta.

Changes in version 1.1b: update for babel-3.6.

Changes in version 1.2: new command \nombre to format numbers; removed command \fup borrowed from the french package (\up does a better job in L^{A}TEX 2ε); also removed aliases \french and \english (frenchb.cfg is a better place for these).

20 The Italian language

The file italian.dtx^{16} It defines all the language-specific macros for the Italian language.

   For this language the \clubpenalty, \widowpenalty and \finalhyphendemerits

are set to rather high values.

21 The Portuguese language

The file portuges.dtx^{17} defines all the language-specific macros for the Portuguese language as well as for the Brasilian version of this language.

For this language the character " is made active. In table 5 an overview is given of its purpose.

"| disable ligature at this position.

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word. "" like "-, but producing no hyphen sign (for words that should break at some sign such as “entrada/salida.”

"< for French left double quotes (similar to <<).
"> for French right double quotes (similar to >>).

\- like the old \-, but allowing hyphenation in the rest of the word.

Table 5: The extra definitions made by portuges.ldf

22 The Spanish language

The file spanish.dtx^{18} defines all the language definition macro’s for the Spanish^{19} language.

16The file described in this section has version number v1.2j and was last revised on 1996/12/29.

The original author is Maurizio Codogno, (mau@beatles.cselt.stet.it).

17The file described in this section has version number v1.2j and was last revised on 1996/12/23. Contributions were made by Jose Pedro Ramalhete (JRAMALHE@CERNVM or Jose-Pedro Ramalhete@MACMAIL) and Arnaldo Viegas de Lima arnaldo@VNET.IBM.COM.

18The file described in this section has version number v3.4g and was last revised on 1997/01/15. The original author is Julio Sánchez, (jsanchez@gmv.es).

19Catalan used to be part of this file but is now on its own file.

This file^{20} incorporates the result of discussions held in the Spanish-TEX^{21} electronic mail list.

For this language the characters ' ~ and " are made active. In table 6 an overview is given of their purpose. These active accent characters behave according

'a an accent that allows hyphenation. Valid for all vow-

els uppercase and lowercase. 'n a n with a tilde. This is included to improve compatibility with FTC. Works for uppercase too.

"| disable ligature at this position.

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word. "" like "-, but producing no hyphen sign (for words that should break at some sign such as “entrada/salida.” \- like the old \-, but allowing hyphenation in the rest

of the word.

"u a u with dieresis allowing hyphenation.
"a feminine ordinal as in 1^{\underline{a}}.
"o masculine ordinal as in 1^{\underline{o}}.
"< for French left double quotes (similar to <<).
"> for French right double quotes (similar to >>).
~n a n with tilde. Works for uppercase too.

Table 6: The extra definitions made by spanish.ldf

to their original definitions if not followed by one of the characters indicated in that table.

This option includes support for working with extended, 8-bit fonts, if available. Old versions of this file based this support on the existance of special macros with names as in Ferguson’s ML-TEX. This is no longer the case. Support is now based on providing an appropriate definition for the accent macros on entry to the Spanish language. This is automatically done by L^{A}TEX 2ε or NFSS2. If T1 encoding is chosen, and provided that adequate hyphenation patterns^{22} exist, it is possible to get better hyphenation for Spanish than before. The easiest way to use the new encoding with L^{A}TEX 2ε to load the package t1enc with \usepackage. This must be done before loading babel.

20In writing this file, many ideas and actual coding solutions have been taken from a number of sources. The language definition files dutch.sty and germanb.sty are the main contributors and are not explicitly mentioned in the sequel. J. L. Braams and Bernd Raichle have given helpful advice. Another source of inspiration is the experience gained in the use of FTC, a software package written by José A. Mañas. The members of the Spanish-TEX list have helped clarify a number of issues. Other sources are explicitly acknowledged when used. If you think that you contributed something and you are not mentioned, please let me (jsanchez@gmv.es) know. I humbly apologize for any omission.

21spanish-tex@goya.eunet.es, subscription requests can be sent to the address listserv@goya.eunet.es. This list is devoted to discussions on support in TEX for Spanish.

Comments on this language option are welcome there or directly to jsanchez@gmv.es.

22One source for such patterns is the archive at ftp.eunet.es that can be accessed by anonymous FTP or electronic mail to ftpmail@goya.eunet.es. They are in the info directory src/TeX/spanish. The list of Frequently Asked Questions with Answers about TEX for Spanish is kept there as well. That list is meant to be a summary of the discussions held in the Spanish-TEX mail list. Warning: It is in Spanish.

If the combination of keyboard and TEX version that the user has is able to produce the accented characters in the T1 enconding, the user could see the accented characters in the editor, greatly improving the readability of the document source. As of today, this is not a recommended method for producing documents for distribution, although it is possible to mechanically translate the document so that the receiver can make use of it. If care is taken to define the encoding needed by the document, the results are pretty portable.

This option file will automatically detect if the T1 encoding is being used and behave appropriately. If any other encoding is being used, the accent macros will be redefined to allow hyphenation on the accented words.

23 The Catalan language

The file catalan.dtx^{23} defines all the language-specific macro’s for the Catalan language.

For this language only the double quote character (") is made active by default. In table 7 an overview is given of the new macros defined and the new meanings of ". Additionally to that, the user can explicitly activate the acute accent or apostrophe (') and/or the grave accent (`) characters by using the activeacute and activegrave options. In that case, the definitions shown in table 8 also become available^{24}.

\lgem geminated-l digraph (similar to l·l). \Lgem produces

the uppercase version. \up Macro to help typing raised ordinals, like 1er. Takes

one argument. \- like the old \-, but allowing hyphenation in the rest

of the word. "i i with diaeresis, allowing hyphenation in the rest of

the word. Valid for the following vowels: i, u (both lowercase and uppercase).

"c c-cedilla (ç). Valid for both uppercase and lowercase

c. "l geminated-l digraph (similar to l·l). Valid for both

uppercase and lowercase l.

"<  French left double quotes (similar to <<).
">  French right double quotes (similar to >>).

"- explicit hyphen sign, allowing hyphenation in the rest

of the word.

"|  disable ligature at this position.

Table 7: Extra definitions made by file catalan.ldf (activated by default)

These active accents characters behave according to their original definitions if not followed by one of the characters indicated in that table.

23The file described in this section has version number v2.2h and was last revised on 1997/01/08.

24Please note that if the acute accent character is active, it is necessary to take special care of coding apostrophes in a way which cannot be confounded with accents. Therefore, it is necessary to type l'{}estri instead of l'estri.

'e acute accented a, allowing hyphenation in the rest of the word. Valid for the following vowels: e, i, o, u (both lowercase and uppercase).

`a grave accented a, allowing hyphenation in the rest of the word. Valid for the following vowels: a, e, o (both lowercase and uppercase).

Table 8: Extra definitions made by file catalan.ldf (activated only when using the options activeacute and activegrave)

24 The Galician language

The file galician.dtx^{25} defines all the language definition macros for the Galician language.

For this language the characters ' ~ and " are made active. In table 9 an overview is given of their purpose. These active accents character behave according

"| disable ligature at this position.

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word. \- like the old \-, but allowing hyphenation in the rest

of the word. 'a an accent that allows hyphenation. Valid for all vow-

els uppercase and lowercase. 'n a n with a tilde. This is included to improve com-

patibility with FTC. Works for uppercase too.

"u a u with dieresis allowing hyphenation.
"a feminine ordinal as in 1^{\underline{a}}.
"o masculine ordinal as in 1^{\underline{o}}.
~n a n with tilde. Works for uppercase too.

Table 9: The extra definitions made by galician.ldf

to their original definitions if not followed by one of the characters indicated in that table.

25 The Romanian language

The file romanian.dtx^{26} defines all the language-specific macros for the Romanian language.

For this language currently no special definitions are needed or available.

25The file described in this section has version number v1.2g and was last revised on 1997/01/15.

26The file described in this section has version number v1.2h and was last revised on 1996/12/23. A contribution was made by Umstatter Horst (hhu@cernvm.cern.ch).

26 The Danish language

The file danish.dtx^{27} defines all the language definition macros for the Danish language.

For this language the character " is made active. In table 10 an overview is given of its purpose.

"| disable ligature at this position.

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word. "" like "-, but producing no hyphen sign (for words that should break at some sign such as “entrada/salida.”

"` lowered double left quotes (looks like ,,)
"' normal double right quotes
"< for French left double quotes (similar to <<).
"> for French right double quotes (similar to >>).

Table 10: The extra definitions made by danish.ldf

27 The Norwegian language

The file norsk.dtx^{28} defines all the language definition macros for the Norwegian language as well as for a new spelling variant ‘nynorsk’ for this language.

For this language currently no special definitions are needed or available.

28 The Swedish language

The file swedish.dtx^{29} defines all the language-specific macros for the Swedish language.

For this language the character " is made active. In table 11 an overview is given of its purpose. The vertical placement of the ”umlaut” in some letters can be controlled this way.

29 The Finnish language

The file finnish.dtx^{30} defines all the language definition macros for the Finnish language.

27The file described in this section has version number v1.3j and was last revised on 1996/12/23.

A contribution was made by Henning Larsen (larsen@cernvm.cern.ch)

28The file described in this section has version number v1.2h and was last revised on 1996/12/23. Contributions were made by Haavard Helstrup (HAAVARD@CERNVM) and Alv Kjetil Holme (HOLMEA@CERNVM); the ‘nynorsk’ variant has been supplied by Per Steinar Iversen iversen@vxcern.cern.ch) and Terje Engeset Petterst (TERJEEP@VSFYS1.FI.UIB.NO).

29The file described in this section has version number v2.1 and was last revised on 1996/12/23. Contributions were made by Sten Hellman (HELLMAN@CERNVM.CERN.CH) and Erik Öshols (erik@ine.kfk.de).

30The file described in this section has version number v1.3j and was last revised on 1997/01/23. A contribution was made by Mikko KANERVA (KANERVA@CERNVM) and Keranen Reino (KERANEN@CERNVM).

"a   \"a, also implemented for A, o and O.
"w gives å, also works for uppercase letters.

"ff for ff to be hyphenated as ff-f, this is also imple-

mented for b, d, f, g, l, m, n, p, r, s, and t.

"| disable ligature at this position.

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word. "" like "-, but producing no hyphen sign (for compound

words with hyphen, e.g. x-""y).

"~ for a compound word mark without a breakpoint.

"= for a compound word mark with a breakpoint, allowing hyphenation in the composing words.

Table 11: The extra definitions made by swedish.sty

For this language the character " is made active. In table 12 an overview is given of its purpose.

"| disable ligature at this position.

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word. "= an explicit hyphen sign for expressions such as

“pakastekaapit ja -arkut”. "" like "-, but producing no hyphen sign (for words that

should break at some sign such as “entrada/salida.”

"` lowered double left quotes (looks like ,,)
"' normal double right quotes
"< for French left double quotes (similar to <<).
"> for French right double quotes (similar to >>).

\- like the old \-, but allowing hyphenation in the rest of the word.

Table 12: The extra definitions made by finnish.ldf

30 The Hungarian language

The file option magyar.dtx^{31} defines all the language definition macros for the Hungarian language.

\ontoday For this language currently the only special definition that is added is the \ontoday command which works like \today but produces a slightly different date format used in expressions suh as ‘on february 10th’.

31The file described in this section has version number v1.3h and was last revised on 1996/12/23. A contribution was made by Attila Koppanyi (attila@cernvm.cern.ch).

Later updates and suggestions by Árpád Bíró (JZP1104@HUSZEG11.bitnet), Istvan Hamecz (hami@ursus.bke.hu) and Horvath Dezso (horvath@pisa.infn.it).

31 The Estonian language

The file estonian.dtx^{32} defines the language definition macro’s for the Estonian language.

This file was written as part of the TWGML project, and borrows heavily from the babel German and Spanish language files germanb.ldf and spanish.ldf.

Estonian has the same umlauts as German (ä, ö, ü), but in addition to this, we have also õ, and two recent characters š and ž, so we need at least two active characters. We shall use " and ~ to type Estonian accents on ASCII keyboards (in the 7-bit character world). Their use is given in table 13. These active accent

~o  \~o, (and uppercase);
"a  \"a, (and uppercase);
"o  \"o, (and uppercase);
"u  \"u, (and uppercase);
~s  \v s, (and uppercase);
~z  \v z, (and uppercase);
"| disable ligature at this position;

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word; \- like the old \-, but allowing hyphenation in the rest

of the word; "` for Estonian low left double quotes (same as Ger-

man);

"' for Estonian right double quotes;
"< for French left double quotes (also rather popular)
"> for French right double quotes.

Table 13: The extra definitions made by estonian.ldf

characters behave according to their original definitions if not followed by one of the characters indicated in that table; the original quote character can be typed using the macro \dq.

We support also the T1 output encoding (and Cork-encoded text input). You can choose the T1 encoding by the command \usepackage[T1]{fontenc}. This package must be loaded before babel. As the standard Estonian hyphenation file eehyph.tex is in the Cork encoding, choosing this encoding will give you better hyphenation.

As mentioned in the Spanish style file, it may happen that some packages fail (usually in a \message). In this case you should change the order of the \usepackage declarations or the order of the style options in \documentclass.

32The file described in this section has version number v1.0e and was last revised on 1996/12/23. The original author is Enn Saar, (saar@aai.ee).

32 The Croatian language

The file croatian.dtx^{33} defines all the language definition macros for the Croatian language.

For this language currently no special definitions are needed or available.

33 The Czech language

The file czech.dtx^{34} defines all the language definition macros for the Czech language.

For this language \frenchspacing is set and two macros \q and \w for easy access to two accents are defined.

The command \q is used with the letters (t, d, l, and L) and adds a ' to them to simulate a ‘hook’ that should be there. The result looks like t’. The command \w is used to put the ring-accent which appears in ångstrøm over the letters u and

U.

34 The Polish language

The file polish.dtx^{35} defines all the language-specific macros for the Polish language.

For this language the character " is made active. In table 14 an overview is given of its purpose.

35 The Slovak language

The file slovak.dtx^{36} defines all the language-specific macros for the Slovak language.

For this language the macro \q is defined. It is used with the letters (t, d, l, and L) and adds a ' to them to simulate a ‘hook’ that should be there. The result looks like t’.

36 The Slovenian language

The file slovene.dtx^{37} defines all the language-specific macros for the Slovenian language.

33The file described in this section has version number v1.3g and was last revised on 1996/12/23. A contribution was made by Alan Paić (paica@cernvm.cern.ch).

34The file described in this section has version number v1.3h and was last revised on 1996/12/23. Contributions were made by Milos Lokajicek (LOKAJICK@CERNVM).

35The file described in this section has version number v1.2d and was last revised on 1996/12/23.

36The file described in this section has version number v1.2i and was last revised on 1996/12/23.

It was written by Jana Chlebikova (chlebik@euromath.dk).

37The file described in this section has version number v1.2i and was last revised on 1996/12/23.

Contributions were made by Danilo Zavrtanik, University of Ljubljana (YU) and Leon Žlajpah (leon.zlajpah@ijs.si).

"a or \aob, for tailed-a (like a̧)
"A or \Aob, for tailed-A (like A̧)
"e or \eob, for tailed-e (like ȩ)
"E or \Eob, for tailed-E (like Ȩ)

"c or \'c, for accented c (like ć), same with uppercase

letters and n,o,s

"l or \lpb{}, for l with stroke (like ł)
"L or \Lpb{}, for L with stroke (like Ł)
"r or \zkb{}, for pointed z (like ż), cf. pronounciation
"R or \Zkb{}, for pointed Z (like Ż)
"z or \'z, for accented z
"Z or \'Z, for accented Z
"| disable ligature at this position.

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word. "" like "-, but producing no hyphen sign (for compund

words with hyphen, e.g. x-""y).

"` for German left double quotes (looks like ,,).
"' for German right double quotes.
"< for French left double quotes (similar to <<).
"> for French right double quotes (similar to >>).

Table 14: The extra definitions made by polish.sty

For this language the character " is made active. In table 15 an overview is given of its purpose. One of the reasons for this is that in the Slovene language some special characters are used.

"c \"c, also implemented for the lowercase and upper-

case s and z. "- an explicit hyphen sign, allowing hyphenation in the

rest of the word. "" like "-, but producing no hyphen sign (for compund

words with hyphen, e.g. x-""y).

"` for Slovene left double quotes (looks like ,,).
"' for Slovene right double quotes.
"< for French left double quotes (similar to <<).
"> for French right double quotes (similar to >>).

Table 15: The extra definitions made by slovene.ldf

37 The Russian language

The file russianb.dtx^{38} defines all the language-specific macros for the Russian language. It needs the file cyrcod for success documentation with Russian encodings (see below).

38The file described in this section has version number ? and was last revised on ?. This file was initially derived from the original version of german.sty, which has some definitions for Russian.

For this language the character " is made active. In table 16 an overview is given of its purpose.

"|  disable ligature at this position.

"- an explicit hyphen sign, allowing hyphenation in the

rest of the word. "" like "-, but producing no hyphen sign (for compund

words with hyphen, e.g. x-""y or some other signs as “disable/enable”).

"~  for a compound word mark without a breakpoint.

"= for a compound word mark with a breakpoint, allow-

ing hyphenation in the composing words.

"`  for German left double quotes (looks like ,,).
"'  for German right double quotes (looks like “).
” "< for French left double quotes (looks like <<).
">  for French right double quotes (looks like >>).

Table 16: The extra definitions made by russianb

The quotes in table 16 can also be typeset by using the commands in table 17.

\glqq  for German left double quotes (looks like ,,).
\grqq  for German right double quotes (looks like “).
” \flqq for French left double quotes (looks like <<).
\frqq  for French right double quotes (looks like >>).
\dq  the original quotes character (").

Table 17: More commands which produce quotes, defined by babel

The quotation marks traditionally used in Russian language were borrowed from other languages (e.g. French and German) so they keep their original names.

38 The Lower Sorbian language

The file lsorbian.dtx^{39} It defines all the language-specific macros for Lower Sorbian.

39 The Upper Sorbian language

The file usorbian.dtx^{40} It defines all the language-specific macros for Upper Sorbian.

Later the definitions from russian.sty version 1.0b (for L^{A}TEX 2.09), russian.sty version v2.5c (for L^{A}TEX2e) and francais.sty version 4.5c and germanb.sty version 2.5c were added.

39The file described in this section has version number v1.0d and was last revised on 1996/12/23. It was written by Eduard Werner (edi@kaihh.hanse.de).

40The file described in this section has version number v1.0e and was last revised on 1996/12/23. It was written by Eduard Werner (edi@kaihh.hanse.de).

40 The Turkish language

The file turkish.dtx^{41} defines all the language definition macros for the Turkish language^{42}.

Turkish typographic rules specify that a little ‘white space’ should be added before the characters ‘:’, ‘!’ and ‘=’. In order to insert this white space automatically these characters are made ‘active’. Also \frenhspacing is set.

41 The Bahasa language

The file bahasa.dtx^{43} defines all the language definition macros for the bahasa indonesia / bahasa melayu language. Bahasa just means ‘language’ in bahasa indonesia / bahasa melayu. Since both national versions of the language use the same writing, although differing in pronounciation, this file can be used for both languages.

For this language currently no special definitions are needed or available.

41The file described in this section has version number v1.2i and was last revised on 1996/12/23.

42Mustafa Burc, z6001@rziris01.rrz.uni-hamburg.de provided the code for this file. It is based on the work by Pierre Mackay

43The file described in this section has version number v1.0e and was last revised on 1996/12/23.